a. Find the area of a trapezoid with bases of 10 inches and 14 inches, and a height of 5 inches.
Answer by venugopalramana(3286)

SEE THE FOLLOWING EXAMPLE FOR DIFFERENT SIZES 9,7 AND 3..... AND DO AS PER THAT
FORMULA FOR AREA OF A TRAPEZIUM IS
A=(1/2)*{SUM OF PARALLEL SIDES}*HEIGHT..........HERE WE HAVE
A=(1/2)(9+7)*3=24 SQ.CM.